# Trajectory Prediction for Autonomous Driving: Progress, Limitations, and Future Directions

### v2

@@ 2.2.3 HD Maps
Latest datasets supplement collected data with High Definition (HD) maps, which provide detailed representations of static road features, such as lane markers, lane edges, pedestrian crossings, road layouts, and stop signs. Typically, the map is presented as a single large file and requires pre-processing to extract locally relevant environmental information. A Region of Interest (ROI), usually encompassing 50–100 meters around the vehicle’s position, is extracted to capture nearby lanes, crosswalks, traffic signs, and other relevant features. Some SOTA datasets, such as NuScenes [66] and Argoverse [67, 68], provide SDKs to query map data and efficiently extract ROIs. This ROI is then transformed into a BEV image, where semantic elements like drivable areas and lane boundaries are encoded using distinct colors or layers. Prediction models can process this BEV image independently, combining it with past trajectories for improved trajectory forecasting [69, 70, 64], or use it in conjunction with sensor data to create a combined rasterized representation [71, 1] (please refer to Figure 5 for illustration).
Incorporating information about the driving environment from HD maps significantly enhances the accuracy of prediction models by constraining predictions to plausible outcomes. However, this approach requires precise localization of the ego-vehicle and is sensitive to localization errors. Additionally, constructing HD maps involves substantial costs and ongoing efforts to update them. To alleviate the human labor associated with constructing HD maps, the research community is actively exploring methods for automatically constructing them from raw sensor data [72, 73, 74, 75, 76]. Combining HD maps with sensor data offers a promising solution to mitigate localization errors and handle changes in the road environment, such as road closures or modifications to traffic signs. However, this approach requires additional computational pre-processing, which impacts the inference time of the prediction pipeline.
@@ 2.3 Modelling Features
Modeling features in trajectory prediction can be broadly categorized into static and dynamic components. Static features refer to the physical layout of the environment, such as road geometry, sidewalks, or walkable areas. Models that include this information can incorporate spatial constraints, e.g., road boundaries, lanes, or walkable regions, into the forecasting process, allowing predicted trajectories to better align with the underlying scene layout. Dynamic features include features related to the target agents and their interactions. One such feature is the agent’s past motion, where the number of past steps is a design choice and varies across models. For example, physics-based models typically rely on the last one or two steps, while deep learning models with attention mechanisms can learn temporal dependencies from longer sequences. Interaction modeling can be beneficial in the scenarios of dense traffic and intersections, where agent behaviors are tightly coupled. In the following sections, we elaborate on each category and describe how different models employ these features.
@@ 2.3.1 Environment Features
Traffic agents rely on the environment’s topology to plan their potential routes, and this geometric information can significantly enhance prediction accuracy [77, 78, 79, 80, 81]. One approach to integrating environmental features is through explicit modelling. Ma and et al. [29] incorporated the distance to the boundaries of crossroads into their prediction model, while Yang and et al. [82] included the presence of traffic lights, stop signs, and zebra crossings. Similarly, Pool et al. [80] incorporated static features for cyclist path prediction by estimating the distance of cyclists to the nearest intersection. This estimation was passed as a contextual feature to a GRU-based RNN, enabling the model to account for the impact of intersections on trajectory predictions.
Rehder et al. [81] presented a more advanced approach by constructing a topological map that organizes semantic features, such as roads, sidewalks, and obstacles, into an occupancy grid. A recurrent Mixture Density Network (rMDN) predicts potential pedestrian destinations, which are integrated into the topological map to encode spatial relationships and constraints. This map serves as the basis for trajectory prediction, ensuring paths align with environmental features.
Another method is implicit modelling, where environmental features are integrated indirectly, focusing on their influence on agent behavior without explicitly extracting or representing them. Bera et al. [83] employs implicit modelling through the Reciprocal Velocity Obstacle (RVO) framework, which adjusts pedestrian trajectories to account for static obstacles like buildings or parked cars without explicitly modelling these elements.
@@ 2.3.2 Target Agents
As discussed in Section 2.2, it is common for prediction methods to generate forecasts based on the past trajectories of target agents [84, 85, 86, 87], and/or incorporate velocity history [88, 41, 89], as well as heading and orientation [90, 91]. In heterogeneous models, including the category of the object (e.g., person, truck, cyclist) enables prediction models [92, 93] to learn category-specific dynamics, although these methods may be prone to errors caused by the incorrect target agent classification by the detector. For vehicles, learning individual behavioral patterns and motion dynamics from historical data [94] can improve prediction accuracy by capturing unique driving styles. Similarly, in pedestrian trajectory prediction, appearance and posture can offer insights into pedestrians’ intentions and planned trajectories. Studies [95, 96, 97] have focused on extracting visual cues to capture behavioral patterns and improve prediction accuracy. To estimate posture, two main approaches are employed: using skeleton key-points [50, 98, 99, 100] or leveraging pretrained CNNs [101, 102, 98].
Incorporating additional semantic attributes requires either integrating a subnetwork within the model to extract these attributes from the input modalities or utilizing networks pre-trained for these tasks, e.g., estimating head direction and gestures in pedestrian behavior prediction. In the first case, the availability of annotated data containing such attributes is often limited, and errors in predicting these attributes can impact overall accuracy. In the second case, the trade-off between the accuracy boost from including these attributes and the added time complexity to the inference pipeline must be carefully evaluated.
@@ 2.3.3 Interaction Features
The integration of autonomous vehicles into human-driven traffic requires a deep understanding of socially compatible communication among traffic agents and the ability to interpret these interactions. To ensure smooth navigation and safe arrival at destinations, trajectory prediction models must account for interaction-related factors [103, 104, 105, 106]. This is particularly critical in dense, interaction-intensive environments, where avoiding collisions depends on accurately modelling the interplay between traffic agents. A simple approach to integrating interactions involves using hand-crafted features, such as relative position and speed of neighboring agents, the distance between vehicles and pedestrians, and time to collision [29, 107, 82, 108]. While straightforward, this method often fails to capture the complexity and variability of interaction patterns in real-world scenarios. More sophisticated techniques for modelling interactions include game theory, social field models, and graph-based approaches.
This section briefly elaborates on Computational Cognition and Social Pooling methods, while GNN-based modelling of interactions between traffic agents is addressed in Section 3.4. Social Pooling is a widely adopted learning-based method in the literature. For a comprehensive review of other approaches and a detailed study on interaction modelling for autonomous driving, readers are referred to the most extensive review available on this topic [109].
Computational cognition methods are based on the premise that interactions between agents are underpinned by cognitive mechanisms, suggesting that a deeper understanding of cognitive science can guide the development of socially-compatible autonomous driving systems [110]. A key theory in this context is Theory of Mind (ToM), which posits that humans can infer the intentions of others by reasoning from their perspective. Cuzzolin et al. [111] argue that integrating ToM into intelligent vehicles is essential for enabling them to share the road with humans effectively, as naive pattern recognition approaches often fail to capture the nuances of spontaneous human behavior. Leveraging knowledge from traffic psychology, particularly about humans’ ability to infer others’ beliefs and intents without explicit communication, can formalize machine ToM for autonomous vehicles [112].
@@ 2.4 Modelling Approach
The choice of a modelling approach depends on the application requirements, platform characteristics, and environmental complexity, such as crowded scenarios or off-road autonomy [120, 121]. Traditional model-based methods, like the Constant-Velocity model or Interactive Multiple Model, explicitly encode motion constraints and are efficient for short-horizon predictions. Physics-based methods are discussed in more details in Section 3.2. For long-horizon predictions, data-driven approaches, which can incorporate spatial and temporal contexts, are more suitable. One of the categories are sampling methods [122, 123, 124, 125, 126, 127, 128, 129], which generate plausible trajectories by modelling the distribution of future trajectories conditioned on the past. These methods estimate the parameters of the distribution from observed data and sample future trajectories given the past trajectory. Depending on the modelling approach, sampling can be parametric, such as Gaussian or Gaussian Mixture Models, or learning-based, like GANs or Autoencoders. For parametric methods, Makansi et al. [122] introduced a two-stage framework that employs a modified Evolving Winner-Takes-All (EWTA) loss during sampling to produce diverse and representative hypotheses from Gaussian distributions. A second network in the framework is used to refine the drawn samples to better align them with the true data distribution. In learning-based methods [123, 124, 125, 126, 127, 128, 129], models are trained to map the relationship between past and future trajectories into a latent space. At inference, given a latent variable and past trajectory, the model generates plausible forecasts. Ma et al. [125] proposed a likelihood-based approach that optimizes for high-likelihood predictions while incorporating diversity-aware losses to represent multiple future behaviors. Gilles et al. [126] introduced a heatmap-driven method that encodes spatial reasoning and agent interactions, producing interpretable likelihood maps for socially-aware multi-agent sampling. We elaborate further on learning-based methods with latent spaces in Sections 3.4.3 and 3.4.4.
@@
On more recent approaches, diffusion models [130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140] have been actively utilized in trajectory forecasting. These models follow a two-phase process: first, Gaussian noise is added to the data to produce a latent representation aligned with a simple prior; then, a reverse process iteratively denoises it, conditioned on past trajectories and scene context. During inference, starting from random noise, the model refines predictions into diverse, multimodal, and context-aware trajectories. Jiang et al. [134] proposed MotionDiffuser, which leverages a permutation-invariant transformer-based denoiser with cross-attention mechanisms to integrate scene context, past trajectories, and agent interactions. It employs PCA to compress trajectories into a latent space for computational efficiency and supports controlled sampling via differentiable cost functions, enabling constraints such as collision avoidance. Barquero et al. [135] introduced the BeLFusion model, a diffusion framework for behavior-driven human motion prediction, which uses a behavior coupler to conditionally sample from a behavioral latent space, ensuring realistic and behaviorally coherent trajectories. Yuan et al. [133] developed PhysDiff, integrating a motion projection module with a physics simulator to iteratively guide the denoising process, eliminating artifacts like floating and ground penetration, and producing physically accurate trajectories.
@@
Continuing with data-driven models, learning-based prediction methods [141, 142] learn motion dynamics from driving data and are well-suited for complex scenarios. However, their generalization and robustness heavily depend on data variability, often requiring additional mechanisms that increase architectural complexity and computational demands. Among Machine Learning (ML) methods, techniques Gaussian Processes, Hidden Markov Models, Dynamic Bayesian Networks, and Gaussian Mixture Models are discussed in Section 3.3. While ML methods are effective in simpler environments with limited interactions, Deep Learning (DL) methods excel at long-term predictions in dynamic environments due to their ability to learn hierarchical and non-linear representations. Section 3.4 discusses key DL approaches, including Sequential Models, Convolutional Neural Networks, Graph Neural Networks, Transformers, Autoencoders, and Generative Adversarial Networks. Another approach, Reinforcement Learning (RL), assumes agents aim to achieve goals by optimizing policies that maximize rewards or minimize costs. RL is particularly suited for trajectory prediction tasks involving interactions and decision-making, as it estimates underlying cost functions to identify optimal policies. RL approaches are further elaborated in Section 3.5.
@@
Modelling interactions between road agents has significantly enhanced the performance of learning-based methods. Interaction-aware approaches [143, 144, 145, 146, 147, 148, 149] explicitly model each agent’s surroundings to capture inter-dependencies between their actions while adhering to social norms. In mixed traffic scenarios, where autonomous vehicles operate alongside human-driven vehicles, achieving safe and efficient navigation requires understanding and interpreting the implicit and explicit ways road users communicate. Incorporating interaction-related factors into trajectory prediction models is essential for enabling autonomous vehicles to avoid collisions, particularly in dense and interaction-intensive environments. This survey discusses interaction modelling across various approaches, emphasizing its importance in developing socially-compatible and reliable trajectory prediction systems. Another line of approaches focuses on intention-aware models. These methods [150, 151, 152, 153, 154, 155, 156] incorporate the vehicle’s maneuver intentions to predict its future states more accurately. Maneuvers are defined as short-term, goal-oriented decisions comprising a sequence of continuous states that collectively lead to achieving a global objective. They can be broadly categorized into two types: lateral maneuvers (e.g., maintaining the current lane, turning left or right, changing lanes) and longitudinal maneuvers (e.g., maintaining speed, accelerating, or decelerating). Integrating intention-awareness introduces an intermediate layer into prediction models, where future forecasts are informed by planning-based reasoning rather than being treated as random or purely reactive paths.
@@ 2.5 Output Modalities
The primary classification of trajectory prediction outputs is based on whether they are unimodal or multimodal. Unimodal outputs are deterministic trajectory points or parameters of a unimodal trajectory distribution. Several studies [22, 92, 27, 24] focus on generating unimodal predictions. However, a significant limitation of unimodal approaches is their inability to capture the inherently multimodal nature of driving behaviour. For instance, in a lane with arrows indicating both straight and left turns, a vehicle could either proceed straight or turn left. Until the driver signals their intent (e.g., by activating an indicator or beginning to incline left), both options should be considered as equally plausible. Ignoring these multimodal characteristics makes it difficult to ensure collision avoidance in complex traffic scenarios, such as lane merging or navigating intersections. Multimodal outputs address this limitation by generating predictions as a multimodal probability distribution [19, 20, 64]. These outputs better align with real-world traffic scenarios, though they present greater challenges in training. A key advantage of multimodal distributions is their ability to associate probabilities with each possible trajectory.
@@
A portion of studies focuses on predicting maneuver intentions as output [18, 17, 21, 26, 60]. Common maneuver categories include changing lanes, staying in the current lane, going straight, and turning left or right. For the joint prediction of trajectories and maneuvers, two main approaches exist. The first approach involves multi-task models [157], where maneuvers intentions and trajectories are predicted as separate outputs but share a common feature extraction module. The second approach [158, 159, 160, 101, 102] predicts maneuvers and trajectories separately, with the predicted maneuvers used to guide trajectory prediction.
@@
Lastly, occupancy grid and occupancy flow are alternative output formats for prediction models. These methods [161, 162, 163] estimate the future occupancy of each cell in a BEV map of the driving area. Occupancy grids provide a spatial representation by predicting the probability of each grid cell being occupied at future time steps [164, 165] and directly estimating spatial probabilities for the entire scene. However, the accuracy of these predictions is constrained by the resolution of the grid cells. A potential solution is to use a flexible grid resolution that adapts to the environment: finer grids for dense urban traffic and coarser grids for highways, optimizing the balance between precision and computational efficiency. Occupancy flow extends the concept of occupancy grids by not only estimating the future occupancy of grid cells but also capturing the motion dynamics within these cells [166, 167, 168]. While occupancy grids encode only spatial information, occupancy flow incorporates temporal changes by predicting both the future occupancy and the flow, reflecting how the objects will move. By modelling motion dynamics at the grid-cell level, occupancy flow eliminates the need for tracker, as it encodes both spatial and temporal information directly from the input data. Both formats are well suited for complex traffic scenarios and smooth integration with planning module.

### v3

@@ 2.2.3 HD Maps
Latest datasets supplement collected data with High Definition (HD) maps, which provide detailed representations of static road features, such as lane markers, lane edges, pedestrian crossings, road layouts, and stop signs. Typically, the map is presented as a single large file and requires pre-processing to extract locally relevant environmental information. A Region of Interest (ROI), usually encompassing 50–100 meters around the vehicle’s position, is extracted to capture nearby lanes, crosswalks, traffic signs, and other relevant features. Some SOTA datasets, such as NuScenes [17] and Argoverse [68, 69], provide SDKs to query map data and efficiently extract ROIs. This ROI is then transformed into a BEV image, where semantic elements like drivable areas and lane boundaries are encoded using distinct colors or layers. Prediction models can process this BEV image independently, combining it with past trajectories for improved trajectory forecasting [70, 71, 66], or use it in conjunction with sensor data to create a combined rasterized representation [72, 1] (refer to Figure 5 for illustration).
Incorporating information about the driving environment from HD maps and geometrically constraining future trajectories can boost forecasting accuracy in structured environments such as urban roads, intersections, and highways. Regarding large empirical drops in displacement error with map information: VectorNet [73] achieves a 28% reduction at a 3-second horizon by incorporating lane polylines, and LaneGCN [74] reduces 6-second error from 1.90 to 1.35 meters by encoding lane centerlines. More moderate gains are reported by the model QCNet [75], where the addition of a map-based attention mechanism contributed to a reduction in displacement error from 0.82 to 0.78 meters. The effectiveness of integrating map data depends on how it is represented — whether as vectorized lane graphs, local semantic priors, or attention-weighted sequences, and how it is integrated with agent dynamics.
As for limitations, this approach requires precise localization of the ego-vehicle and is sensitive to localization errors. Additionally, constructing HD maps involves substantial costs and ongoing efforts to update them. To alleviate the human labor associated with constructing HD maps, the research community is actively exploring methods for automatically constructing them from raw sensor data [76, 77, 78, 79, 80]. Combining HD maps with sensor data offers a promising solution to mitigate localization errors and handle changes in the road environment, such as road closures or modifications to traffic signs, however, it requires additional computational pre-processing, which impacts the inference time of the prediction pipeline.
@@ 2.3 Modelling Features
Modeling features in trajectory prediction can be broadly categorized into static and dynamic components. Static features refer to the physical layout of the environment, such as road geometry, sidewalks, or walkable areas. Models that include this information can incorporate spatial constraints, e.g., road boundaries, lanes, or walkable regions, into the forecasting process, allowing predicted trajectories to better align with the underlying scene layout. Dynamic features include features related to the target agents and their interactions. Agent’s past motion is a dynamic feature, where the number of past steps is a design choice and varies across models. For example, physics-based models typically rely on the last one or two steps, while deep learning models with attention mechanisms can learn temporal dependencies from longer sequences. Interaction modeling can be beneficial in the scenarios of dense traffic and intersections, where agent behaviors are tightly coupled. In the following sections, we elaborate on each category and describe how different models employ these features.
@@ 2.3.1 Environment Features
Traffic agents rely on the environment’s topology to plan their potential routes, and this geometric information can be integrated to enhance prediction accuracy [81, 82, 83, 84, 85]. The environmental features can be integrated through explicit modelling. Ma and et al. [31] incorporated the distance to the boundaries of crossroads into their prediction model, while Yang and et al. [86] included the presence of traffic lights, stop signs, and zebra crossings. Similarly, Pool et al. [84] incorporated static features for cyclist path prediction by estimating the distance of cyclists to the nearest intersection. This estimation was passed as a contextual feature to a GRU-based RNN, enabling the model to account for the impact of intersections on trajectory predictions. Rehder et al. [85] presented a more advanced approach by constructing a topological map that organizes semantic features, such as roads, sidewalks, and obstacles, into an occupancy grid. A recurrent Mixture Density Network (rMDN) predicts potential pedestrian destinations, which are integrated into the topological map to encode spatial relationships and constraints.
@@ 2.3.2 Target Agents
As discussed in Section 2.2, it is common for prediction methods to generate forecasts based on the past trajectories of target agents [87, 88, 89, 90], and/or incorporate velocity history [91, 43, 92], as well as heading and orientation [93, 94]. In heterogeneous models, including the category of the object (e.g., person, truck, cyclist) enables prediction models [95, 96] to learn category-specific dynamics, although these methods may be prone to errors caused by the incorrect target agent classification by the detector. For vehicles, learning individual behavioral patterns historical data [97] can improve prediction accuracy by capturing unique driving styles. Similarly, in pedestrian trajectory prediction, appearance cues and posture can offer insights into pedestrians’ intentions. Studies [98, 99, 100] have focused on extracting visual cues to capture behavioral patterns to improve prediction accuracy. To estimate posture, two main approaches are employed: using skeleton key-points [52, 101, 102, 103] or leveraging pretrained CNNs [104, 105, 101].
Incorporating additional semantic attributes requires either integrating a subnetwork within the model to extract these attributes from the input modalities or utilizing networks pre-trained for these tasks, e.g., estimating head direction and gestures in pedestrian behavior prediction. In the first case, the availability of annotated data containing such attributes is often limited, and errors in predicting these attributes can propagate. In the second case, the trade-off between the accuracy boost from including these attributes and the added time complexity to the inference pipeline needs to be evaluated while model testing.
@@ 2.3.3 Interaction Features
The integration of autonomous vehicles into human-driven traffic requires AV to understand socially compatible communication among traffic agents and the ability to interpret these interactions. Accounting for interaction-related factors [106, 107, 108, 109] is beneficial in dense, interaction-intensive environments, where avoiding collisions depends on accurately modelling the interplay between traffic agents. A simple approach to integrating interactions involves using hand-crafted features, such as relative position and speed of neighboring agents, the distance between vehicles and pedestrians, and time to collision [31, 110, 86, 111]. While straightforward, this method often fails to capture the complexity and variability of interaction patterns in real-world scenarios. More sophisticated techniques for modelling interactions include social field models, game theory and graph-based approaches and computation cognition methods [112, 113, 114].
Further, we briefly elaborate on Social Pooling methods, which are widely adopted learning-based method in the literature. GNN-based modelling of interactions between traffic agents is discussed in Section 3.4. For a comprehensive review of other approaches and a detailed study on interaction modelling for autonomous driving, readers are referred to the extensive review available on this topic [115].
@@
The performance of interaction-aware models is heavily influenced by how interaction learning is integrated, as well as by the diversity of scenarios in the dataset (only a subset of which typically involves densely interactive settings). Quantitatively, AgentFormer [123], a transformer-based model that jointly encodes social and temporal cues, demonstrates through ablations that modeling either cue in isolation leads to degraded performance. Qualitatively, models like Social-LSTM [116] and Social-GAN [122] have shown improvements in generating more socially acceptable trajectories in crowded environments. Regarding the efficiency of social pooling, models such as Social-STGCNN [124] and Social-BiGAT [125] by replacing pooling operations with graph convolutions or attention mechanisms, enhanced predictive accuracy and reduced inference latency.
@@ 2.4 Modelling Approach
The choice of a modelling approach depends on the application requirements, platform characteristics, and environmental complexity, e.g., crowded urban scenarios or off-road autonomy [126, 127]. Traditional model-based methods, such as the Constant-Velocity model or Interactive Multiple Model, explicitly encode motion constraints to generate physically plausible trajectories. These methods are simple in terms of implementation and interpretation. A more detailed discussion is provided in Section 3.2. To capture complex trajectory patterns over longer horizons, the community often adopts data-driven approaches capable of capturing both spatial and temporal contexts. One such category is sampling-based methods [128, 129, 130, 131, 132, 133, 134, 135], which first model the distribution of future trajectories conditioned on observed past paths and then generate samples from this distribution to produce future predictions. Sampling can be parametric, such as Gaussian or Gaussian Mixture Models, or learning-based, like GANs or Autoencoders. For parametric methods, Makansi et al. [128] introduced a two-stage framework that employs a modified Evolving Winner-Takes-All (EWTA) loss during sampling to produce diverse and representative hypotheses from Gaussian distributions. A second network in the framework is used to refine the drawn samples to better align them with the true data distribution. In learning-based methods [129, 130, 131, 132, 133, 134, 135], models are trained to map the relationship between past and future trajectories into a latent space. At inference, given a latent variable and past trajectory, the model generates plausible forecasts. Gilles et al. [132] introduced a heatmap-driven method that encodes spatial reasoning and agent interactions, producing interpretable likelihood maps for socially-aware multi-agent sampling. We elaborate further on learning-based methods with latent spaces in Sections 3.4.3 and 3.4.4.
@@
On more recent approaches, diffusion models [136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146] have been actively utilized in trajectory forecasting. These models follow a two-phase process: first, Gaussian noise is added to the data to produce a latent representation aligned with a prior; then, a reverse process iteratively denoises it, conditioned on past trajectories and scene context. During inference, starting from random noise, the model refines predictions into diverse, multimodal, and context-aware trajectories. Jiang et al. [140] proposed MotionDiffuser, which leverages a permutation-invariant transformer-based denoiser with cross-attention mechanisms to integrate scene context, past trajectories, and agent interactions. It employs PCA to compress trajectories into a latent space for computational efficiency and supports controlled sampling via differentiable cost functions, enabling constraints such as collision avoidance. Barquero et al. [141] introduced the BeLFusion model, a diffusion framework for human motion prediction, which uses a behavior coupler to conditionally sample from a behavioral latent space, ensuring realistic and behaviorally coherent trajectories. Yuan et al. [139] developed PhysDiff, integrating a motion projection module with a physics simulator to iteratively guide the denoising process, eliminating artifacts like floating and ground penetration, and producing physically accurate trajectories.
@@
Continuing with data-driven models, learning-based prediction methods [147, 148] aim to learn motion dynamics directly from driving data. In classical Machine Learning (ML) approaches such as Gaussian Processes, Hidden Markov Models, Dynamic Bayesian Networks, and Gaussian Mixture Models (see Section 3.3), model parameters, including distribution parameters, transition probabilities, and kernel hyperparameters (for GP) — are estimated from data using probabilistic inference. While ML methods offer interpretability grounded in probabilistic reasoning and perform well in simpler environments, Deep Learning (DL) methods excel in complex, dynamic settings due to their ability to learn hierarchical and non-linear representations. Nevertheless, their generalization is highly sensitive to the diversity and quality of training data, often requiring architectural enhancements or training strategies to improve robustness, which can increase computational cost. Section 3.4 discusses key DL approaches, including Sequential Models, Convolutional Neural Networks, Graph Neural Networks, Transformers, Autoencoders, and Generative Adversarial Networks. The mentioned methods are reviewed in this paper within the supervised learning setting. Self-supervised learning (SSL) approaches [149, 150, 151] in trajectory forecasting are commonly employed to learn transferable representations of motion patterns, addressing the scarcity of labeled datasets by leveraging large volumes of unlabeled data. SSL is typically used to enhance the quality of learned features, assisting the prediction model in better generalization. A common setup involves multi-task training, where SSL objectives are combined with supervised forecasting tasks. These methods often rely on proxy tasks, which do not require manual labels but encourage the model to learn meaningful structure from the data, e.g., in contrastive learning, the model is trained to bring representations of similar trajectories closer together while pushing apart dissimilar ones, thereby encouraging the encoding of discriminative motion features. The specifics of SSL methods are out of the scope of this review, we refer the reader to the study [152] for more details. Another approach, Reinforcement Learning (RL), assumes agents aim to achieve goals by optimizing policies that maximize rewards or minimize costs. RL is particularly suited for trajectory prediction tasks involving interactions and decision-making, as it estimates underlying cost functions to identify optimal policies. RL approaches are further elaborated in Section 3.5.
@@
In mixed traffic, where autonomous vehicles operate alongside human-driven vehicles, safe navigation requires understanding and interpreting the implicit and explicit ways road users interact. Incorporating interaction-related factors into trajectory prediction models [153, 154, 155, 156, 157, 158, 159] can improve AV operation in dense, crowded environments. This survey discusses interaction modeling across various approaches, highlighting its benefits in developing socially compatible trajectory prediction systems. Another line of approaches focuses on intention-aware models. These methods [160, 161, 162, 163, 164, 165, 166] incorporate the vehicle’s maneuver intentions to predict its future states. Maneuvers are defined as short-term, goal-oriented decisions comprising a sequence of continuous states that collectively lead to achieving a global objective. They can be broadly categorized into two types: lateral maneuvers (e.g., maintaining the current lane, turning left or right, changing lanes) and longitudinal maneuvers (e.g., maintaining speed, accelerating, or decelerating). Integrating intention-awareness introduces an intermediate layer into prediction models, where future forecasts are informed by planning-based reasoning.
@@ 2.5 Output Modalities
For object-centric trajectory prediction, where forecasts are centered around observed agents, prediction outputs are categorized as unimodal and multimodal. Unimodal outputs are either deterministic trajectory points or parameters of a unimodal trajectory distribution [24, 95, 29, 26]. However, driving behavior is inherently multimodal. For instance, in a lane with arrows indicating both straight and left turns, a vehicle could either proceed straight or turn left. Until the driver signals their intent (e.g., by activating an indicator or beginning to incline left), both options should be considered equally plausible. Ignoring such multimodal characteristics can negatively affect tasks like collision avoidance at intersections or during lane merging. Therefore, models with multimodal outputs [21, 22, 66] generate predictions as a multimodal probability distribution to better reflect real-world traffic behavior. These models can produce a set of plausible future trajectories, each associated with a probability that reflects the likelihood of that path being followed.
@@
A set of studies focuses on predicting maneuver intentions as the output [20, 19, 23, 28, 62]. Common maneuver categories include changing lanes, staying in the current lane, going straight, and turning left or right. For the joint prediction of trajectories and maneuvers, two main approaches exist. The first approach involves multi-task models [167], where maneuvers intentions and trajectories are predicted as separate outputs but share a common feature extraction module. The second approach [168, 169, 170, 104, 105] predicts maneuvers and trajectories separately, with the predicted maneuvers used to guide trajectory prediction.
@@
For occupancy-centric models, the environment is typically divided into a grid by a perception module, and the prediction module [171, 172, 173] estimates the likelihood of each cell being occupied at future time steps [174, 175]. Occupancy grids are very natural for integration with the planning module; however, the resolution of the grid controls the balance between precision and computational demands. A potential solution is to use a flexible grid resolution that adapts to the environment: finer grids for dense urban traffic and coarser grids for highways. Occupancy flow extends the concept of occupancy grids by also capturing the motion dynamics within these cells [176, 177, 178]. While occupancy grids generate a sequence of static occupancy maps for future time steps, occupancy flow explicitly models dynamics by jointly predicting occupancy and flow fields, encoding how occupied regions are expected to move. Both formats learn from sequences of input frames and do not require tracking, as they derive spatial and temporal information directly from the input data.